# Geometry-Aware Preference Learning for 3D Texture Generation

### v1

@@ 2.3.2 Geometry-Guided Texture Colorization Reward
Note that to compute the curvature-quided colorization reward, we need the mean curvature 2D map as a way to incorporate mesh’s geometry information into texture learning objectives (Equation 4). To this end, we construct a mean curvature 2D map, a 2D image in UV space that encodes the mean curvature of the underlying 3D mesh surface and then compute the rewards in the 2D UV space by comparing the texure image and the 2D curvature map. This process involves the following key steps.
@@
2.3.3 Texture Features Emphasis Reward
This reward encourages texture patterns that emphasize 3D surface structure while maintaining perceptual richness through color variation. It consists of two components: (i) texture-curvature magnitude alignment which strengthens texture variations in high-curvature regions to enhance capturing geometric structures, and (ii) colorfulness inspired by (Du et al., 2024; Hasler & Suesstrunk, 2003), which discourages desaturated textures by encouraging vibrant color use, measured through the standard deviation and mean distance of opponent color channels from neutral gray. Let $T(x,y)\in[0,1]^{3}$ denote the RGB texture at pixel $(x,y)$ and $C(x,y)\in[0,1]$ be the normalized curvature map in UV space. The texture-curvature alignment term is mathematically defined as the negative mean squared error (MSE):
where $\nabla\mathrm{T}(x,y)=\sqrt{\left\|\frac{\partial\mathrm{T}}{\partial x}(x,y)\right\|^{2}+\left\|\frac{\partial\mathrm{T}}{\partial y}(x,y)\right\|^{2}}$ is the texture gradient magnitude and $N$ is the total number of pixels in the texture image. Then, we augment this reward with a colorfulness term based on opponent color channels (Hasler & Suesstrunk, 2003). Let $R(x,y),G(x,y),B(x,y)\in\mathbb{R}^{H\times W}$ denote the red, green, and blue channels of the texture. We compute the opponent axes:
@@ 2.3.3 Texture Features Emphasis Reward
$R_{3}=\alpha_{m}R_{magnitude}+\alpha_{c}R_{color}$ (9)
2.3.4 Symmetry-Aware Texture Generation Reward
We propose a differentiable reward function that encourages texture patterns to be symmetric across mirrored UV coordinates while maintaining visual richness through color variation. We first identify the mesh’s intrinsic plane of symmetry using Principal Component Analysis (PCA) (Abdi & Williams, 2010; Pearson, 1901), then compute mirror pairs of surface points across this plane, project these pairs into the 2D UV domain for use in a symmetry-aware reward function, and finally compute the symmetry reward. The four steps are outlined in the following.
Symmetry plane estimation via PCA. To estimate the mesh’s dominant symmetry plane, we apply principal component analysis (PCA) to the 3D vertext coordinates. Let $\{{v_{i}}\in\mathbb{R}^{3}\}_{i=1}^{N}$ denoted the 3D vertext positions of the mesh. We then compute the centroid of the mesh as $\mathbf{c}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{v}_{i}$. Next, we construct the covariance matrix of the centered vertices: $\Sigma=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbf{v}_{i}-\mathbf{c}\right)\left(\mathbf{v}_{i}-\mathbf{c}\right)^{\top}$. We then perform eigen-decomposition of $\Sigma$, yielding eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq$ and corresponding orthonormal eigenvectors $e_{1},e_{2},e_{3}$. Teh eigenvector $n=e_{1}$ associated with teh smallest eigenvalue is to taken to define the normal vector of the symmetry plane, as it corresponds to the direction of least variance in the point cloud (vertices). The symmetry plane $\Pi$ is then defined as the set of points:
Reflecting vertices across the symmetry plane. To identify symmetric correspondences, we first compute the signed distance of each vertex to the plane: $d_{i}=\left(\mathbf{v}_{i}-\mathbf{c}\right)\cdot\mathbf{n}$. We keep only the vertices where $d_{i}>0$, i.e., those lying on one side of the plane, to avoid duplication. Each selected $v_{i}$ is then reflected across the plane to compute its mirrored position $v_{i}^{mirror}\in\mathbb{R}^{3}$ using: $\mathbf{v}_{i}^{\text{mirror }}=\mathbf{v}_{i}-2\left[\left(\mathbf{v}_{i}-\mathbf{c}\right)\cdot\mathbf{n}\right]\mathbf{n}$.
Projecting 3D mirror points into UV space. To obtain the UV coordinates corresponding to $v_{i}^{mirror}$, we perform a closest-point query on the mesh surface to find the nearest triangle containing the surface projection of $v_{i}^{mirror}$. Given the triangle’s 3D vertices $v_{1},v_{2},v_{3}$, and corresponding UV coordinates $u_{1},u_{2},u_{3}\in[0,1]^{2}$, we compute barycentric coordinates $(\beta_{1},\beta_{2},\beta_{3})$ of the projected point and obtain the UV coordinates of the mirror as: $\mathbf{u}_{i}^{\text{mirror }}=\beta_{1}\mathbf{u}_{1}+\beta_{2}\mathbf{u}_{2}+\beta_{3}\mathbf{u}_{3}$. We store the UV coordinate $u_{i}$ of the original vertex along with $u_{i}^{mirror}$ and form a pair $(u_{i},u_{i}^{mirror})$ used in our symmetry-awared reward.
Symmetry-aware texture reward design Given a set of UV coordinate pairs $\{(\mathbf{u}_{i},\mathbf{u}_{i}^{\text{mirror}})\}_{i=1}^{M}$, obtained by projecting symmetric 3D vertex pairs into the UV domain, we define a differentiable reward function that encourages symmetry in the generated texture. Let $\mathcal{T}\in\mathbb{R}^{3\times H\times W}$ denote the RGB texture defined in UV space. For each UV pair, we sample RGB values via bilinear interpolation:
where $\mathcal{T}_{i},\mathcal{T}_{i}^{\text{mirror}}\in\mathbb{R}^{3}$ represent the RGB values at each location. Sampling is implemented using the differentiable grid_sample operation in Pytorch (Paszke, 2019) to ensure gradient flow during training. To encourage symmetry, we minimize the mean squared error (MSE) between original and mirrored samples:
We define the symmetry reward as the negative of this loss:
$\mathcal{R}_{\text{symmetry}}=-\mathcal{L}_{\text{symmetry}}$ (13)
To discourage trivial (e.g., grayscale) solutions, similar to texture features emphasis reward (Section 2.3.3), we augment the symmetry reward with a colorfulness term using Equation 8. The total reward is a weighted combination of the symmetry alignment and colorfulness terms:
Figure 2: Qualitative results of the geometry-texture alignment experiment on a rabbit (bunny) mesh. For each example, we show the rendered 3D object from different single viewpoints. Minimum curvature vectors, representing the underlying surface geometry, are visualized in red color in the last two columns overlaid on the textured objects for comparison. As shown, our method produces textures whose patterns align more closely with the mesh’s curvature directions, unlike InTex (Tang et al., 2024). Moreover, a notable outcome in our results is the emergence of repetitive texture patterns after fine-tuning with the geometry-texture alignment reward. This behavior arises from the differentiable sampling strategy used during reward computation. Specifically, it encourages the model to place edge features at specific UV coordinates which ultimately results in structured and repeated patterns in the texture (see the last part of Figure 5)
3 Experiments and Results
Table 1: A quantitative comparison between our proposed method (after fine-tuning) and InTeX (Tang et al., 2024) (before fine-tuning) across four experiments. We repeat each experiment five times and report the mean and standard deviation values of rewards. GEO-TEX-ALIGN, GEO-TEX-COLOR, TEX-EMPHASIS, and SYM-AWARE stand for geometry-texture alignment reward (Section 2.3.1), geometry-guided texture colorization reward (Section 2.3.2) , texture features emphasis reward (Section 2.3.3), and symmetry-aware texture generation reward (Section 2.3.4), respectively.
Reward Ours InTeX
Geo-Tex-Align 0.3347 $\pm$ 0.0039 0.2915 $\pm$ 0.0019
Geo-Tex-Color 0.6006 $\pm$ 0.0003 -0.0256 $\pm$ 0.0195
Tex-Emphasis -0.1106 $\pm$ 0.0002 -0.1423 $\pm$ 0.0007
Sym-Aware 0.0114 $\pm$ 0.0093 -0.0635 $\pm$ 0.0039
To show the effectiveness of our approach, we perform four training experiments using the four proposed geomatry-aware reward functions in Section 2.3, each detailed in the following. We qualitatively and quantitatively compare our results against the the automated version of InTeX’s (Tang et al., 2024) method, an state-of-the-art method in the field of texture generation. We report the qualitative results for each experiment in Figure 2, Figure 3, Figure 4, and Figure 5 and quantitative results in Table 1 (more results in Appendix D).
Geometry-Texture Alignment. The objective of this task is to generate texture images whose patterns are aligned with the surface geometry of a 3D mesh object, represented by its minimal curvature directions. Figure 2 presents qualitative results on a rabbit mesh textured using different textual prompts. Each row corresponds to a different description, with the main texture pattern written on the left and overlaid vectors indicating the minimal curvature directions on the surface. As illustrated, our method successfully guides texture patterns to follow the underlying geometric cues, a capability that InTex (Tang et al., 2024) lacks (more results in Figure 10). Compared to InTeX, which uses a pre-trained model without fine-tuning, our approach exhibits an interesting capability: the generation of repetitive texture patterns that align with the curvature structure. This behavior arises from the differentiable sampling step we introduce during the geometry-texture alignment stage, where texture gradients are sampled at the same spatial locations as curvature vectors. By consistently associating these positions with edge-like features in the texture, the model learns to reinforce structure-aware repetition. Additional analysis of this effect are provided in the last part of Section 3.
Geometry-Guided Texture Colorization. This task aims to colorize textures based on surface bending intensity, represented by mean curvature, an average of the minimal and maximal curvature directions, on the 3D mesh. Specifically, the model is encouraged to apply warm colors (e.g., red, yellow) in high-curvature regions and cool colors (e.g., blue, green) in low-curvature areas. Figure 3 shows qualitative results on rabbit and cow mesh objects colorirzed with different textual prompts. As illustrated, our method consistently adapts texture colors, regardless of initial patterns, according to local curvature and successfully maps warmth and coolness to geometric variation (more results in Figure 11).
Figure 3: Qualitative results of the geometry-guided texture colorization experiment on a cow mesh object. We show the rendered 3D cow from multiple viewpoints. The goal is to colorize textures based on surface bending intensity, represented by mean curvature, an average of the minimal and maximal curvature directions, on the 3D mesh. Specifically, the model is encouraged to apply warm colors (e.g. yellow) in high-curvature regions and cool colors (e.g., blue) in low-curvature areas. As illustrated, our method consistently adapts texture colors, regardless of initial patterns, according to local curvature and successfully maps warmth and coolness to geometric variation.
Texrure Features Emphasis. This task aims to learn texture images with salient features (e.g., edges) emphasized at regions of high surface bending, represented by the magnitude of mean curvature. This promotes texture patterns that highlight 3D surface structure while preserving perceptual richness through color variation. Figure 4 shows qualitative results on a rabbit mesh with brick patterns. As illustrated, our method enhances texture features, such as edges and mortar, in proportion to local curvature, a capability InTex (Tang et al., 2024) lacks, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit (more results in Figure 12).
Figure 4: Qualitative results of the texture features emphasis experiment on a rabbit (bunny) object. We show the rendered 3D bunny from multiple viewpoints. This goal is to learn texture images with salient features (e.g., edges) emphasized at regions of high surface bending, represented by mean curvature. This encourages texture patterns that highlight 3D surface structure while preserving perceptual richness through color variation. As illustrated, our method enhances texture features, such as edges and mortar, in proportion to mean curvature, a capability InTex (Tang et al., 2024) lacks, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit.
Symmetry-Aware Texture Generation. The goal of this experiment is to encourages texture consistency across symmetric regions of a 3D object (see Section 2.3.4). We evaluate this by training our pipeline using the symmetry reward on a balloon mesh object. Figure 5 presents qualitative results before and after fine-tuning with the symmetry reward. We show the rendered 3D object from multiple viewpoints, alongside the corresponding texture images (the last row), which highlight the symmetric regions. A vertical dashed line marks the symmetry axis in each texture image. The purple plane passing through the center of the balloon in each viewpoint indicates the estimated symmetry plane of the object. As shown, compared to the pre-trained model (Tang et al., 2024), our method generates textures that are more consistent across symmetric parts of the mesh, demonstrating the reward’s effectiveness in enforcing symmetry consistency. Without symmetry supervision, patterns often differ noticeably between sides (more results in Figure 9).
Figure 5: Qualitative results of the symmetry-aware experiment on a 3D balloon mesh object. We show the rendered 3D object from multiple viewpoints, alongside the corresponding texture images (the last row), which highlight the symmetric regions. A vertical dashed line marks the symmetry axis in each texture image. The purple plane passing through the center of the balloon in each viewpoint indicates the estimated symmetry plane of the object. As shown, compared to the pre-trained model (Tang et al., 2024), our method generates textures that are more consistent across symmetric parts of the mesh, demonstrating the reward’s effectiveness in enforcing symmetry consistency. Without symmetry supervision, patterns often differ noticeably between sides.
Emergence of Repetitive Patterns through Geometry-Aligned Sampling. An interesting observation in our results (see Figure 2 and Figure 10) is the emergence of repetitive texture patterns after fine-tuning with the geometry-texture alignment reward. This behavior stems from the mismatch between the number of curvature vectors and the number of texture gradients. Specifically, curvature is a vertex-level attribute defined per mesh vertex, whereas texture gradients are defined per pixel in the 2D UV space. For example, in the bunny mesh example, we have 4487 curvature vectors (one per vertex), but over one million texture gradients from a $1024\times 1024$ texture image. To compute cosine similarity between curvature vectors and texture gradients, both must be in the same space and of the same dimension. We address this by sampling the texture gradients at the UV coordinates corresponding to mesh vertices, the same positions where curvature vectors are defined. This results in two aligned tensors (both of size 4487 in the bunny case), enabling the reward function to compare directional alignment directly. Over time, this differentiable sampling mechanism guides the model to place strong texture features (e.g., edges) at those specific UV coordinates. As these positions remain fixed (being tied to the mesh vertices), the model learns to reinforce edge features at the same UV locations across views and iterations. This alignment results in visually repetitive patterns in the generated textures, especially in regions of high curvature.
4 Conclusions
This study presents an end-to-end differentiable preference learning framework for 3D texture generation, enabling human preferences, expressed as differentiable reward functions, to be back-propagated through the entire 3D generative pipeline. This makes the process inherently geometry-aware. To demonstrate its effectiveness, we introduce four geometry-aware reward functions that guide texture generation to align closely with the input mesh geometry. The framework is broadly applicable, supporting any differentiable feedback defined in either 2D texture space or directly on the 3D surface, and is extensible to tasks beyond texture generation. As future work, we envision applying this approach to general 3D content creation, including joint optimization of geometry and texture.
Acknowledgements
We acknowledge funding from Mila Tech Transfer Grant with Silicolabs. Compute resources provided by Calcul Quebec and the digital researcher alliance.
Appendix A Remarks on The Texture Generation Step
Due to the stochastic nature of the diffusion models, applying the iterative painting process for texturing process, explained in Section 2.1, would leads to having inconsistent textures with noticeable seams on the appearance of the object. To resolve this issue, we take a dynamic partitioning approach similar to (Richardson et al., 2023; Chen et al., 2023), where each rendered viewpoint is divided to three regions based on a concept called view direction cosine. The view direction cosine is utilized to determine how the surface is facing the viewer. Mathematically, the view direction cosine is represented as the cosine of the angle between two vectors: the view direction vector which is the direction of the viewer’s (in our case, the camera) line of sight, and mesh’s surface normal vectors associated with each face or vertex of the mesh. These normals indicate the direction perpendicular to the surface at that point (vertex):
[…]
where $\vec{V}$ is the view vector, a vector starting from the camera position points toward a point on the surface of the 3D mesh object, $\vec{N}$ is the surface normal, vector perpendicular to the mesh surface, and $\phi$ is the angle between the view vector and surface normal. During the rendering process, this helps determine how to project and render 3D objects onto a 2D screen, ensuring that objects are viewed from the correct perspective. More specifically, the view direction cosine is computed for each rendered viewpoint, and the rendered viewpoint is partitioned into three regions based on the value of view direction cosine: 1) generate - the regions have been viewed for the first time by the camera and have never been painted during the texturing process, 2) refine - the regions that have already been viewed and painted from a viewpoint in previous iterations, but is now seen from a better camera angle (i.e. higher view direction cosine values) and should be painted again, and 3) keep - the regions that have already been viewed and pained from a good camera angle and should not be painted again. Section A.1 presents further details and the differentiable mathematical representation of our texture generation step.
A.1 Mathematical Representation
Mathematically, the texture-generation process can be formulated as steps below (see Figure 6): (i) rendering, render the object using a differentiable renderer (Laine et al., 2020) and extract a rendering buffer $h^{i}$ corresponding to the camera viewpoint $i$, including painted viewpoint image, depth maps, normal maps, and UV coordinates:
$h^{i}=f_{render}(a,v_{gen})$ (16)
where $a$ is the 3D asset’s parameters including texture image, the view direction cosine for the current view, view direction cosine for previous iterations, $v_{gen}$ is the camera viewpoints for texture-generation process, and $i\in v_{gen}$ is the current viewpoint. Using the normal map, obtained from the differentiable renderer, we then compute the view direction cosine, and then generate three regions (masks), explained above, for each viewpoint: $M_{generate},M_{refine}$ and $M_{keep}$. These three regions will serve as input to the next step of the texturing process to enforce the consistency in the output texture. Figure 7 visually demonstrates the details of the output generated by the differentiable renderer. (ii) Depth-aware painting, at each denoising step $t$, the depth-aware diffusion process can be formulated as below:
where $f_{controlnet}$ represents the ControlNet model whose parameters are fixed during the training, $h^{i}$ is the differentiable renderer output, $y^{i}_{t-1}$ is the control signal generated by the ControlNet model at timestep $t$ for viewpoint $i$, and $c$ is the prompt embdedding obtained from the CLIP encoder model (Radford et al., 2021). To enforce the consistency during the painting process, we follow (Avrahami et al., 2023) to incorporate latent mask blending in the latent space of the diffusion process. Specifically, by referring to the part that we wish to modify as $m_{blended}$ and to the remaining part as $m_{keep}=1-m_{blended}$, we blend the two parts in the latent space, as the diffusion progresses. This modifies the sampling process in a way that the diffusion model does not change the keep regions of the rendered image. Therefore, the latent variable at the current diffusion timestep $t$ is computed as:
A.2 Modifications To Make Texture Generation Step Differentiable
Differentiable Camera Pose Computation. Instead of using traditional camera positioning, to enable end-to-end training with backpropagation through the camera positioning steps, we provide a re-implementation of this procedure using PyTorch (Paszke, 2019) operations, preserving gradient flow. More specifically, camera poses are computed using statically generated camera-to-world transformation matrices from spherical coordinates (elevation and azimuth). This process involved the following steps: (i) spherical to cartesian conversion - given azimuth $\phi$, elevation $\theta$, and radius $r$, the camera position $c\in\mathbb{R}^{3}$ is computed as:
$x=r\cos(\theta)\sin(\phi),$ (20)
$y=-r\sin(\theta),$
$z=r\cos(\theta)\cos(\phi),$
$\mathbf{c}=[x,y,z]^{\top}+\mathbf{t},$
[…]
Differentiable Viewpoint Dynamic Partitioning (Masking). In the original implementation of the texture generation step (Tang et al., 2024), dynamic viewpoint partitioning for multi-view texturing was performed using hard, non-differentiable binary masks based on thresholding and bitwise operations. Specifically, three disjoint masks were used to guide different behaviors across image regions during training: (i) a generation mask $M_{generate}$, (ii) a refinement mask $M_{refine}$, and (iii) a keep mask $M_{keep}$. Let $c\in\mathbb{R}^{H\times W}$ denotes the count map (indicating coverage across viewpoints), and $v,v_{old}\in\mathbb{R}^{H\times W}$ be the view-consistency scores for the current and cached viewpoints, respectively. The original masks were defined as:
$M_{\text{generate }}$ $=1[c<0.1]$ (23)
$M_{\text{keep }}$ =$=\neg M_{\text{generate }}\cdot\neg M_{\text{refine }}$
where $1[\cdot]$ denotes the indicator function and logical operators like, $\neg$ produce discrete masks, making them non-differentiable and unsuitable for gradient-based optimization. To enable end-to-end differentiability, we replace these hard decisions with soft, continuous approximations using the sigmoid function. Let $\sigma_{k}(x)=\frac{1}{1+\exp(-kx)}$ denote the sigmoid function with steepness $k$. We define soft masks as follows: (i) soft generation mask -
$M_{\text{generate }}^{\text{soft }}=1-\sigma_{k}(c-0.1)$ (24)
Activation Checkpointing. Gradient or activation checkpointing (Chen et al., 2016) is a technique that trades compute for memory. Instead of keeping tensors needed for backward alive until they are used in gradient computation during backward, forward computation in checkpointed regions omits saving tensors for backward and recomputes them during the backward pass. In our framework, we apply two levels of gradient checkpointing to our pipeline. First, low-level checkpointing, where we apply the checkpointing technique to the sampling process in the diffusion model by only storing the input latent for each denoising step, and re-compute the UNet activations during the backpropagation. Second, high-level checkpointing, where we apply the checkpointing technique to every single-view texturing process in our multi-view texturing algorithm. Practically, what it means is that, for each viewpoint during the multi-view iterative texturing process, intermediate tensors inside the per-view texturing are not saved. Instead, only the inputs (camera pose, prompt embeddings, etc.) are saved. Then, during the backward pass, we recompute the forward pass of the per-view texturing again so gradients can flow backward. Overall, this technique allows to keep memory usage constant even across many viewpoints and diffusion steps which consequently leads to scaling to more views and using higher-resolution renders or more diffusion steps to have a more realistic and visually appealing texture image without any memory issues.
Low-Rank Adaptation (LoRA). LoRA is a technique, originally introduced in (Hu et al., 2022) for large language models, that keeps the pre-trained model weights fixed and adds new trainable low-rank matrices into each layer of the neural architecture which highly reduces the number of trainable parameters during the training process. Mathematically, for a layer with base parameters $W$ whose forward pass is $h=Wx$, the LoRA adapted layer is $h=Wx+BAx$, where $A$ and $B$ are the low-rank trainable matrices and the $W$ remains unchanged during the training process. In our case, we apply the LoRA techniques to the U-Net architecture inside the latent diffusion model (LDM) (Liu et al., 2019) which lies at the heart of our texture generation step. This allows us to fine-tune the LDM using a parameter set that is smaller by several orders of magnitude, around 1000 times fewer, than the full set of LDM parameters.
Appendix C Remarks on Principal Curvature Directions
Principal curvature directions at a point on a surface describe how the surface bends in different directions. Curvature is defined as $k=\frac{1}{R}$ where $R$ is the radius of the circle that best fits the curve in a given direction. Among all possible directions, two orthogonal directions, called the principal curvature directions, are of special interest: the maximum and minimum curvature directions. The minimum curvature direction corresponds to the direction in which the surface bends the least (i.e., the largest fitting circle), while the maximum curvature direction corresponds to the direction in which it bends the most (i.e., the smallest fitting circle). See Figure 8 for an illustration. Additionally, the mean curvature at a point is defined as the average of the two principal curvatures: $H=\frac{1}{2}(k_{1}+k_{2})$. We use the Libigl library (Jacobson et al., 2018) to compute both the mean curvature values and the principal curvature directions on our 3D mesh surfaces.
Appendix D Additional Experiments and Results

### v2

@@ F.1 Geometry-Aware Texture Colorization Reward Design
Note that to compute the curvature-quided colorization reward, we need the mean curvature 2D map as a way to incorporate mesh’s geometry information into texture learning objectives (Eq. 27). To this end, we construct a mean curvature 2D map, a 2D image in UV space that encodes the mean curvature of the underlying 3D mesh surface and then compute the rewards in the 2D UV space by comparing the texure image and the 2D curvature map. This process involves the following key steps.
@@
F.2 Geometry-Aware Texture Colorization Experiment and Results
This task aims to colorize textures based on surface bending intensity, represented by mean curvature, an average of the minimal and maximal curvature directions, on the 3D mesh. Specifically, the model is encouraged to apply warm colors (e.g., red, yellow) in high-curvature regions and cool colors (e.g., blue, green) in low-curvature areas. Fig. 11 shows qualitative results on rabbit and cow mesh objects colorirzed with different textual prompts. As illustrated, our method consistently adapts texture colors, regardless of initial patterns, according to local curvature and successfully maps warmth and coolness to geometric variation (more results in Fig. 11).
Appendix G Attempts to Overcome Reward Hacking
To prevent the fine-tuned diffusion model from diverging too far from the original pre-trained model and to mitigate issues such as overfitting and reward hacking, we incorporate three regularization strategies: KL divergence [11], LoRA scaling [31], and a colorfulness incentive [12]. These techniques are added as regularization terms to each reward function we design and are evaluated for their effectiveness during training. While KL regularization offers a principled approach to reducing reward overfitting, we find that LoRA scaling and the colorfulness incentive [12] consistently yield better empirical results. We detail each of these techniques in the following.
KL-Divergence. Inspired by [7] and similar to [11], we introduce a regularization penalty that encourages the fine-tuned model to stay close to the behavior of the pre-trained Stable Diffusion model. Specifically, we penalize the squared difference between the noise predictions of the pre-trained and fine-tuned U-Net models at the final denoising step. Let $\epsilon_{\mathrm{pt}}$ and $\epsilon_{\mathrm{ft}}$ denote the predicted noise from the pre-trained and fine-tuned models, respectively. We define the penalty term as:
$\beta_{\mathrm{KL}}\left\|\epsilon_{\mathrm{ft}}-\epsilon_{\mathrm{pt}}\right\|_{2}^{2}$ (29)
and add it to the overall reward during training. The coefficient $\beta_{\mathrm{KL}}$ controls the strength of this regularization and is varied across different experimental settings. Intuitively, this penalty discourages the fine-tuned model from diverging too far from its pre-trained counterpart, helping to mitigate issues such as catastrophic forgetting or overly narrow specialization. This approach is justified by the fact that, under fixed variance, the KL divergence between two Gaussian noise distributions simplifies to the squared difference between their means. In the context of conditioned diffusion models, noise prediction $\epsilon$ is conditioned on a noised latent $x_{t}$. We compute this penalty at the final denoising timestep $t=1$, just before image reconstruction. For each training batch, we compute $\epsilon_{\mathrm{pt}}$ by forwarding the frozen pre-trained model (with gradients disabled), and $\epsilon_{\mathrm{ft}}$ by forwarding the fine-tuned model. The squared difference, scaled by $\beta_{\mathrm{KL}}$, is then added to the reward signal.
LoRA scaling. Inspired by [7] and similar to [31], we scale the LoRA parameters by a factor in the range (0,1](0,1]( 0 , 1 ] to modulate their effect during fine-tuning. Empirically, we observe that factors close to 1, specifically within [0.9,1][0.9,1][ 0.9 , 1 ], consistently yield the best performance in our experiments.
Colorfulness incentive. Inspired by [10, 12], we discourage desaturated texture images by encouraging vibrant color use, measured through the standard deviation and mean distance of opponent color channels from neutral gray. To this end, we augment our reward with a colorfulness term based on opponent color channels [12]. Let $R(x,y),G(x,y),B(x,y)\in\mathbb{R}^{H\times W}$ denote the red, green, and blue channels of the texture. We compute the opponent axes:
@@ Appendix G Attempts to Overcome Reward Hacking
where $\sigma_{rg}$, $\sigma_{yb}$, $\mu_{rg}$, and $\mu_{yb}$ are the standard deviations and mean colors offsets corresponding to color opponent components $rg(x,y)=R(x,y)-G(x,y)$ and $yb(x,y)=\frac{R(x,y)+G(x,y)}{2}-B(x,y)$, respectively. This term is incorporated into both our texture feature emphasis and symmetry-aware texture generation reward functions to promote visually rich and saturated outputs.
Appendix H Additional Experiments and Results